OpenStudy (melaniemfeith):

A larger number is double the sum of 3 and a smaller number. The larger number is 2 less than 3 times the smaller number. If y represents the larger number and x represents the smaller number, which equations model the situation?

OpenStudy (prathamesh_m):

The sum of 3 and the smaller number would be given by x + 3. The larger number is double this amount so the first equation would be: y = 2(x + 3)

OpenStudy (robtobey2):

\[\{2 (x+3)=y,y=3 x-2\} \]x=8, y=22